OpenStudy (anonymous):

Let y = (7/x^3)(√196x^6 +364x^2 + (169/x^2)) Find the derivative of y. Hint: Simplify y before computing the derivative. Answer: (-364/x^5)

OpenStudy (anonymous):

does anyone know how to simplify the equation first before taking it's derivative?

OpenStudy (anonymous):

can you rewrite the equation using the equation tool on this site. It is tough to see in this form.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[y = \frac{ 7 }{ x^3 }\sqrt{196x^6 + 364x^2 + \frac{ 169 }{ x^2 }}\]

OpenStudy (anonymous):

Answer is \[\frac{ dy }{ dx } = \frac{ -364 }{ x^5 }\]

OpenStudy (anonymous):

I don't know how to simplify the radical. I would only know how to differentiate right away and it'll get messy

OpenStudy (anonymous):

ok thanks anyway

OpenStudy (anonymous):

you would use the product rule right?

OpenStudy (anonymous):

ya.

OpenStudy (anonymous):

I'm doing it too. it's getting pretty rough

OpenStudy (anonymous):

\[y' = -21x ^{-4} (196x^6 + 364x^2 + 169x^{-2})^{1/2} + 7x ^{-3}(\frac{ 1 }{ 2 } (196x^6 + 364 + 169x ^{-2})^{-1/2}(1176x^5 +728x -338x ^{-3})^{-1/2}\]

OpenStudy (anonymous):

whoops lol

OpenStudy (anonymous):

simplifying is the hardest part

zepdrix (zepdrix):

Hmm there is something going on inside this square root, I'm trying to figure it out. Not quite there yet, but I think it's not long off. 196 = 14^2 169=13^2 Hmmmm

zepdrix (zepdrix):

OH! yes found it! :)

zepdrix (zepdrix):

364 = 2 (13*14)

zepdrix (zepdrix):

(14x^4 + 13)^2 I think it'll simplify like that.

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

so how would this look like

OpenStudy (anonymous):

after simplifying?

zepdrix (zepdrix):

 Hmm, I think it simplifies like this, looks much easier to deal with :) I got lazy on the multiplying there hehe